# Mitigating Metropolitan Carbon Emissions with Dynamic Eco-driving at Scale

### v1

@@ 4.1 Reinforcement Learning
$\pi^{*}(s)=\operatorname*{argmax}_{\pi}\mathop{\mathbb{E}}\left[\sum_{t=0}^{H}{\gamma^{t}r(s_{t},a_{t})|s_{0},\pi}\right]$ (1)
@@ 4.2 Contextual Reinforcement Learning
$\pi^{*}(s)=\operatorname*{argmax}_{\pi}\mathop{\mathbb{E}}\left[\sum_{c\in\mathcal{C}}\sum_{t=0}^{H}{\gamma^{t}r_{c}(s_{t},a_{t})|s_{0}^{c},\pi}\right]$ (2)
@@ 4.3 Multi-task Learning
$\pi^{*}(s)=\operatorname*{argmax}_{\pi}\mathop{\mathbb{E}}\left[\sum_{\tau\in\mathcal{T}}\sum_{t=0}^{H}{\gamma^{t}r_{\tau}(s_{t},a_{t})|s_{0}^{\tau},\pi}\right]$ (3)
@@ 4.4 Zero-shot Transfer Learning
Zero-shot transfer learning extends the concept of learning across multiple tasks, as in cMDPs, where the agent must perform well on new, unseen tasks without additional training. The goal is to leverage knowledge from a set of source tasks $\mathcal{T}_{\text{source}}$, which the agent has experienced during training, to generalize effectively to a set of target tasks $\mathcal{T}_{\text{target}}$. These target tasks are typically drawn from a distribution that may encompass tasks not encountered during training. In particular, a policy $\pi_{source}$ trained with the $\mathcal{T}_{\text{source}}$ is evaluated on the target tasks $\mathcal{T}_{\text{target}}$ without further training on the target tasks. Such a transfer learning strategy reduces the training overhead and can be better than training a policy from scratch on the target tasks Kirk et al. .
@@ 5.1 Regional eco-driving problem
Optimizing eco-driving at the traffic network scale can yield the highest benefits, but it is a significantly hard problem and intractable even with advances in modern computing and thus remains an open optimization challenge Qadri et al. (2020). Therefore, many studies adopt a network decomposition from the full network to the intersection level and optimize vehicle trajectories at individual intersections. While this has been a standard practice, it has to be performed carefully as isolated optimization of intersection-level behaviors can cause unexpected adverse events (e.g., intersection queue spillback from one intersection to another) at the network level. In this work, we adopt the same network decomposition strategy and take careful precautions to isolate the effects of the changed vehicle behaviors on each individual intersection.
In light of this, consider a region comprised of traffic scenario variations $\Phi$ where $|\Phi|$ is very large (e.g., millions of individual signalized intersection scenarios). With the network decomposition strategy, traffic scenario variations are stemming at the intersection level (e.g., changes in traffic flow, temperature, etc.). Let $\pi^{e}\coloneqq(\pi^{e}_{\phi})_{\phi\in\Phi}$ denote the set of eco-driving control laws. Similarly, let $\pi^{b}\coloneqq(\pi^{b}_{\phi})_{\phi\in\Phi}$ denote the baseline (status quo driving). Let $f(\pi,\phi)$ denote a function that captures the CO2 emission for the scenario $\phi$ under the control law $\pi$. We then define the regional eco-driving effectiveness as,
$E_{\Phi}(\pi^{e},\pi^{b})\coloneqq 1-\frac{\operatorname*{\mathbb{E}}_{\phi\in\Phi}[f(\pi^{e}_{\phi},\phi)]}{\operatorname*{\mathbb{E}}_{\phi\in\Phi}[f(\pi^{b}_{\phi},\phi)]}$ (4)
We thus seek to solve the regional eco-driving problem for the control laws $\pi^{e}$ such that,
Find
$\pi^{e}=\arg\max_{\pi}E_{\Phi}(\pi,\pi^{b})$ (5)
@@
$n_{\phi}^{e}\geq n_{\phi}^{b}\;\;\;\;\forall\;\phi\in\Phi$ (6)
$m_{\phi}^{e}\leq\tau\;\;\;\;\forall\;\phi\in\Phi$ (7)
$w_{\phi}^{e}\leq w_{\phi}^{b}\;\;\;\;\forall\;\phi\in\Phi$ (8)
Here, the $n_{\phi}^{e}$ and $n_{\phi}^{b}$ denote the throughput of the intersection where the eco-driving is assessed in the scenario $\phi$ under the control laws of $\pi^{e}_{\phi}$ and $\pi^{b}_{\phi}$, respectively. Similarly, $m_{\phi}^{e}$ denotes the average observed queue length to lane length ratio of the immediately following intersection to the intersection where the eco-driving is assessed, and $\tau$ is a hyperparameter. Finally, $w_{\phi}^{e}$ and $w_{\phi}^{b}$ denote the average waiting time per vehicle in the immediate preceding intersection to the intersection where the eco-driving is assessed under control laws of $\pi^{e}_{\phi}$ and $\pi^{b}_{\phi}$, respectively.
The Equations 6, 7, and 8 ensure the network decomposition strategy would not induce traffic events that would adversely affect the traffic flow at the network level. In particular, Equations 6 ensures that by eco-driving, we do not reduce the intersection throughout, as a reduction in throughput could create traffic queues that would adversely affect the network traffic flow. Equations 7 ensure that if there is an increase in the throughput at an intersection with eco-driving, it will not lead to a saturated incoming approach in the immediately following intersection, which ensures that the immediately following intersection has the capacity to absorb the increased throughput. Finally, Equation 8 ensures that if the eco-driving affects the inflow of vehicles to the intersection being assessed, it would not be propagated back to the intersection preceding it. These constraints collectively ensure that changes in the traffic flow due to eco-driving are either localized to the assessed intersection or implemented with high confidence that they will not adversely affect overall network traffic flow. While this conservative approach may yield a lower bound of potential benefits, it provides a reliable assessment of the impact of eco-driving considering the network scale traffic flow dynamics.
@@ 5.2 Eco-driving control problem at an intersection
$a_{i}(t)=\begin{cases}x_{i}(s_{i}(t),\dot{s}_{i}(t),v_{i}(t),p)&\\ \quad\quad\quad\quad\quad\quad\text{if vehicle $i$ is controlled to eco-drive}\\ y_{i}(s_{i}(t),\dot{s}_{i}(t),v_{i}(t),p)&\\ \quad\quad\quad\quad\quad\quad\text{otherwise}\end{cases}$ (10)
@@
Here, $n$ represents the total number of vehicles in the fleet, $a_{i}(t)$, $v_{i}(t)$, $s_{i}(t)$, and $\dot{s}(t)$ denote the acceleration, velocity, headway and relative velocity of vehicle $i$ at time $t$, $E(\cdot)$ denotes the emission function that measures instantaneous emission of vehicle $i$ at time $t$, $T_{i}$ denotes the travel time of vehicle $i$, and $d$ represents the total required travel distance, which includes both arrival to and departure from the intersection. Furthermore, $\Delta t$ defines the discretized time step usually in $[0.5,1]$.
The minimum and maximum headway values are denoted as $s_{min}$ and $s_{max}$, while $v_{min}$, $v_{max}$ and $a_{min}$, $a_{max}$ indicate the same for velocities and accelerations. These variables and constraints ensure safety, connectivity via V2V communication (if applicable), and passenger comfort.
We use IDM Treiber et al. (2000) as the dynamics model for human drivers. Thus, $y_{j}$ entails the ordinary differential equations defined by the calibrated parameterization of IDM for human-driven vehicle $j$. The $x_{i}$ for controlled vehicle $i$ is learned using multi-task deep reinforcement learning.
@@ 6.1 Scenario Modeling
For any given scenario, we capture the emissions effects of these factors through three interrelated levels of modeling: 1) traffic-level, 2) vehicle-level, and 3) behavior-level as shown in Figure 3. All simulation scenario modeling is performed in the open-source agent-based SUMO microscopic traffic simulator Lopez et al. (2018).
@@
In vehicle-level modeling, we aim to capture the factors owing to the operating characteristics of individual vehicle engines. These factors are modeled using vehicle emission models. This approach allows us to account for factors such as temperature, humidity, seasonality, and weather conditions. The industry-standard Motor Vehicle Emission Simulator (MOVES) by the US Environmental Protection Agency epa (a) provides accurate exhaust emissions for this purpose. Yet, it is impractical for our large-scale analysis, where we must assess millions of scenarios, given the MOVES takes about three minutes per scenario. To overcome this, we create a suite of efficient surrogate neural network emission models using MOVES as the data source Sanchez et al. (2022, 2023). Leveraging MOVES also allows us to capture the cumulative average effects of vehicle maintenance, vehicle model, vehicle characteristics, road types, aerodynamic drag, air conditioning, and vehicle load.
Last, we focus on behavior-level modeling, which accounts for the factors owing to altered vehicle behaviors due to eco-driving optimization. These factors are modeled using deep reinforcement learning and rule-based control modeling. We consider behaviors with a substantial impact on emissions, such as idling, speeding, acceleration, deceleration, pulse-and-glide, and lane changes Xu et al. (2021). Behavior-level modeling across diverse traffic scenarios is a core challenge addressed in this work, as discussed next.
@@ 6.2 Eco-driving Control
Eco-driving control involves devising a set of eco-driving strategies for the representative scenario set $\Phi^{{}^{\prime}}$. Eco-driving control for one million scenarios is a challenging task. Traditional methods like dynamic programming are computationally limiting Doan et al. (2018), while heuristics are often suboptimal Katsaros et al. (2011). Trajectory optimizations are limited by the complexity of required vehicle dynamics models, particularly due to inter-vehicle dynamics and environmental variability introduced by many factors we consider Yang et al. (2016).
Meanwhile, recent studies have shown the effectiveness of deep reinforcement learning in complex control tasks Wurman et al. (2022); Silver et al. (2017); Bellemare et al. (2020). Neural vehicle control using deep reinforcement learning Wu et al. (2021); Yan et al. (2022); Jayawardana and Wu (2022), which can handle diverse traffic scenarios without manual fine-tuning, is also gaining popularity. However, training such a policy for each scenario is impractical due to the number of scenarios. Training a unified policy is complex due to multifaceted factors and the induced complex expectation Zhang and Yang (2021). Learning both discrete lane changes and continuous longitudinal accelerations presents optimization challenges, demanding diverse exploration strategies and distinct neural network representations.
@@
Similar scenario partitioning: In this step, our aim is to partition scenarios by their similarities, allowing us to train separate MT-DRL policies for each partition. This strategy is inspired by the mixture of expert architecture widely used in real-world systems with generalization requirements Masoudnia and Ebrahimpour (2014). It can reduce scenario diversity within a partition and help reduce training complexity. While each scenario is defined by many factors, we only utilize intersection geometry and traffic flow-specific features as partitioning features because they strongly influence policy behaviors.
However, since signalized intersections lack rotational invariance, comparing them for similarity is challenging. Therefore, we break down the intersection environment into non-conflicting pairs of incoming and outgoing approaches. This has added advantages. It accelerates training by simulating fewer vehicles and improves training stability by alleviating credit assignment challenges Zhou et al. (2020). Further, it has no impact on modeling realism due to the decoupling nature of traffic signals (e.g., northbound traffic and eastbound traffic do not collide). These selected pairs, along with the ghost approaches connected to them, thus become our training environments and are illustrated in Figure 7.
@@
The training environments are then partitioned into 28 partitions, considering the semi-autonomous vehicle penetration level, lane count, and traffic signal phase count as partition features. However, the environments in each partition may not form a uniform distribution. This may make environments sampled in consecutive training iterations differ significantly, requiring adaptive learning strategies. To overcome this, we adopt a synthetic environment distribution $\Phi_{T}$ derived from the full scenario set $\Phi$ for each partition with a uniform distribution for policy training as illustrated in Figure 2. Additional details about the training environment distributions can be found in Appendix 10.7.
@@
Observation: Vehicle observation design is mainly influenced by the currently available technology, such as the requisite vehicle sensors, and the significance attributed by existing literature to specific information required for eco-driving. Thus, we define the observation for a vehicle (agent) $i$ as follows.
@@
Reward: The reward function must encourage algorithmic generalization within the cMDP while considering various competing objectives like emission reduction and travel time efficiency. Thus, the reward function for each semi-autonomous vehicle is expressed as $r_{t}^{i}=v_{t}^{i}+\alpha\mathbbm{1}_{v^{i}_{t}<\tau}+\beta e_{t}^{i}$, where $v_{t}^{i}$ represents the velocity of vehicle $i$ at time $t$, and $e_{t}^{i}$ represents its CO2 emissions. In this equation, $\alpha$, $\beta$, and $\tau$ are hyperparameters. The indicator function $\mathbbm{1}_{v^{i}_{t}<\tau}$ indicates whether the vehicle is stopped, while the term $e_{t}^{i}$ encourages low emissions. The velocity term captures the effect on travel time.
It’s generally accepted that limiting accelerations and idling reduces emissions Mintsis et al. (2020). But this may create traffic overflows or reduce intersection throughput. To account for these observations, we devise two policies per partition ${\psi_{i}}$: one, denoted as $\pi_{\psi_{i}}^{1}$, penalize idling, while the other, $\pi_{\psi_{i}}^{2}$, places less weight on it. To achieve this we set $\alpha=35$, $\beta=3$ and $\tau=1$ for $\pi_{\psi_{i}}^{1}$ and $\alpha=10$, $\beta=3$ and $\tau=1$ for $\pi_{\psi_{i}}^{2}$.
Hierarchical controller architecture. While our MT-DRL policy controls vehicle accelerations, we leverage a rule-based default SUMO lane-changing controller for lane changes. It captures 1) strategic lane changes, 2) cooperative lane changes, 3) tactical lane changes, and 4) regulatory lane changes Erdmann (2014). At every time step, we first check if the lane-changing controller would predict a lane-changing maneuver for the vehicle. If a lane-changing maneuver is warranted, precedence is given to the lane-changing maneuvers, and the lane-changing controller takes over the control of the vehicle to control both longitudinal and lateral movements of the vehicle for that time step. Otherwise, we use the MT-DRL policy by default to control the longitudinal movements of the vehicle.

### v2

@@ 4 Preliminaries
4.1 Intelligent Driver Model
In our simulations, we use the Intelligent Driver Model (IDM) (48) to simulate the human drivers. IDM defines the instantaneous acceleration $a(t)$ of a vehicle at time $t$. It is given by the Equation 1 where $v_{0},s_{0}$ and $T$ denote desired velocity, space headway, and time headway, respectively. Maximum acceleration is given by $a$, and comfortable braking deceleration is given by $b$. $v(t)$ is the ego-vehicle velocity, $\delta$ is a constant, and $\Delta v(t)$ denotes the velocity difference between the ego vehicle and the vehicle in front of it.
$a(t)=\frac{dv(t)}{dt}=a\left[1-\left(\frac{v(t)}{v_{0}}\right)^{\delta}-\left(\frac{H(v(t),\Delta v(t)}{h(t)}\right)^{2}\right]$ (1)
$H(v(t),\Delta v(t))=s_{0}+\max\left(0,v(t)T+\frac{v(t)\Delta v(t)}{2\sqrt{ab}}\right)$ (2)
@@ 4.2 Reinforcement Learning
$\pi^{*}=\operatorname*{argmax}_{\pi}\mathop{\mathbb{E}}\left[\sum_{t=0}^{H}{\gamma^{t}r(s_{t},a_{t})|s_{0},\pi}\right]$ (3)
@@ 4.3 Contextual Reinforcement Learning
$\pi^{*}=\operatorname*{argmax}_{\pi}\mathop{\mathbb{E}}\left[\sum_{c\in\mathcal{C}}\sum_{t=0}^{H}{\gamma^{t}r_{c}(s_{t},a_{t})|s_{0}^{c},\pi}\right]$ (4)
@@ 4.4 Multi-task Learning
$\pi^{*}=\operatorname*{argmax}_{\pi}\mathop{\mathbb{E}}\left[\sum_{\omega\in\mathcal{T}}\sum_{t=0}^{H}{\gamma^{t}r_{\omega}(s_{t},a_{t})|s_{0}^{\omega},\pi}\right]$ (5)
@@ 4.5 Zero-shot Transfer Learning
Zero-shot transfer learning extends the concept of learning across multiple tasks, as in CMDPs, where the agent must perform well on new, unseen tasks without additional training. The goal is to leverage knowledge from a set of source tasks $\mathcal{T}_{\text{source}}$, which the agent has experienced during training, to generalize effectively to a set of target tasks $\mathcal{T}_{\text{target}}$. These target tasks are typically drawn from a distribution that may encompass tasks not encountered during training. In particular, a policy $\pi$ trained with the $\mathcal{T}_{\text{source}}$ is evaluated on the target tasks $\mathcal{T}_{\text{target}}$ without further training on the target tasks. Such a transfer learning strategy reduces the training overhead and can be more economical than training a policy from scratch on the target tasks (28).
@@ 5.1 Regional eco-driving problem
Optimizing eco-driving at the traffic network scale can yield the highest benefits, but it is a significantly hard problem and intractable even with advances in modern computing, and thus remains an open optimization challenge (39). Therefore, many studies adopt a network decomposition from the full network to the intersection level and optimize vehicle trajectories at individual intersections. While this has been a standard practice, it has to be performed carefully as isolated optimization of intersection-level behaviors can cause adverse events (e.g., intersection queue spillback from one intersection to another) at the network level. In this work, we adopt the same network decomposition strategy and take careful precautions to isolate the effects of the changed vehicle behaviors on each individual intersection.
In light of this, consider a region comprised of traffic scenario variations $\Phi$ where $|\Phi|$ is very large (e.g., millions of individual signalized intersection scenarios). With the network decomposition strategy, the traffic scenario variations are stemming at the intersection level (e.g., changes in traffic flow, temperature, etc.). Let $\pi\coloneqq(\pi_{\phi})_{\phi\in\Phi}$ denotes a set of driving control laws. Similarly, let $\pi^{b}\coloneqq(\pi^{b}_{\phi})_{\phi\in\Phi}$ denotes the baseline (status quo driving). Let $f(\pi^{\prime},\phi)$ denote a function that captures the CO2 emission for the scenario $\phi$ under a control law $\pi^{\prime}\in\{\pi,\pi^{b}\}$. We then define the regional eco-driving effectiveness as follows. Note that each scenario $\phi$ implicitly encapsulates its own dynamics, which are abstracted within $\phi$ in this formulation.
$E_{\Phi}(\pi,\pi^{b})\coloneqq 1-\frac{\operatorname*{\mathbb{E}}_{\phi\in\Phi}[f(\pi_{\phi},\phi)]}{\operatorname*{\mathbb{E}}_{\phi\in\Phi}[f(\pi^{b}_{\phi},\phi)]}$ (6)
We thus seek to solve the regional eco-driving problem for the best eco-driving control laws $\pi^{e}$ while avoiding negative system impacts to throughput, queue length, and waiting time. That is, we find:
$\pi^{e}=\arg\max_{\pi}E_{\Phi}(\pi,\pi^{b})=\arg\min_{\pi}\operatorname*{\mathbb{E}}_{\phi\in\Phi}[f(\pi_{\phi},\phi)]$ (7)
@@
$n(\pi_{\phi},\phi)\geq n(\pi^{b}_{\phi},\phi)\;\;\;\;\forall\;\phi\in\Phi$ (8)
$m(\pi_{\phi},\phi)\leq\tau\;\;\;\;\forall\;\phi\in\Phi$ (9)
$w(\pi_{\phi},\phi)\leq w(\pi^{b}_{\phi},\phi)\;\;\;\;\forall\;\phi\in\Phi$ (10)
Here, $n(\cdot)$ defines the intersection throughput dynamics function that takes the scenario $\phi$ and the control laws $\pi_{\phi}$ and $\pi^{b}_{\phi}$ and outputs the intersection throughput when the given control policy is applied to the scenario. Similarly, $m(\cdot)$ defines a function that captures the average observed queue length to lane length ratio of the immediately following intersection to the intersection where the eco-driving is assessed, and $\tau$ is a hyperparameter. Finally, $w(\cdot)$ defines a function that captures the average waiting time per vehicle in the immediate preceding intersection to the intersection where the eco-driving is assessed under control laws of $\pi_{\phi}$ and $\pi^{b}_{\phi}$. The functions $n(\cdot)$, $m(\cdot)$, and $w(\cdot)$ serve as global abstractions of the underlying physical traffic phenomena and takes in scenario $\phi$ and the relevant control policy to make it scenario specific (e.g., to capture the effects of each scenario dynamics) to capture how vehicle behaviors and interactions evolve under different control laws in a compact and measurable form.
The Equations 8, 9, and 10 ensure the network decomposition strategy would not induce traffic events that would adversely affect the traffic flow at the network level. In particular, Equations 8 ensures that by eco-driving, we do not reduce the intersection throughout, as a reduction in throughput could create traffic queues that would adversely affect the network traffic flow. Equations 9 ensure that if there is an increase in the throughput at an intersection with eco-driving, it will not lead to a saturated incoming approach in the immediately following intersection, which ensures that the immediately following intersection has the capacity to absorb the increased throughput. Finally, Equation 10 ensures that if the eco-driving affects the inflow of vehicles to the intersection being assessed, it would not be propagated back to the intersection preceding it. These constraints ensure that changes in the traffic flow due to eco-driving are either localized to the assessed intersection or implemented with high confidence that they will not adversely affect overall network traffic flow. While this conservative approach may yield a lower bound of potential benefits, it provides a reliable assessment of the impact of eco-driving, considering the network scale traffic flow dynamics.
@@ 5.2 Eco-driving control problem at an intersection
$a_{i}(t)=\begin{cases}x_{i}(s_{i}(t),\dot{s}_{i}(t),v_{i}(t),p)&\quad\text{if vehicle $i$ is controlled to eco-drive}\\ y_{i}(s_{i}(t),\dot{s}_{i}(t),v_{i}(t),p)&\quad\text{otherwise}\end{cases}$ (12)
@@
Here, $n$ represents the total number of vehicles in the fleet, $a_{i}(t)$, $v_{i}(t)$, $s_{i}(t)$, and $\dot{s}(t)$ denote the acceleration, velocity, headway and relative velocity of vehicle $i$ at time $t$, $E(\cdot)$ denotes the emission function that measures instantaneous emission of vehicle $i$ at time $t$, the multi-objective trade-off hyperparameter is $\alpha$, and $T_{i}$ denotes the travel time of vehicle $i$. The vehicle dynamics are coupled through relative position (headway) and relative velocity. The total required travel distance is denoted by $d$, and includes both arrival to and departure from the intersection. Furthermore, $\Delta t$ defines the discretized time step, usually in $[0.5,1]$.
In Eq 11, we aim to capture the total emission objective value of each vehicle (first summation term) over the entire planning horizon of each vehicle (second summation term). This objective has to be achieved given the minimum and maximum headway values denoted as $s_{min}$ and $s_{max}$, while $v_{min}$, $v_{max}$ and $a_{min}$, $a_{max}$ indicating the same for velocities and accelerations. These variables and their corresponding constraints ensure safety, connectivity via V2V communication (if applicable), and passenger comfort.
We use IDM (47) as the dynamics model for human drivers. Thus, $y_{j}(\cdot)$ entails the ordinary differential equations defined by the calibrated parameterization of IDM for human-driven vehicle $j$. The $x_{i}(\cdot)$ for controlled vehicle $i$ is learned using multi-task deep reinforcement learning.
@@ 6.1 Scenario Modeling
For any given scenario, we capture the emissions effects of these factors through three interrelated levels of modeling: 1) traffic-level, 2) vehicle-level, and 3) behavior-level as shown in Figure 3. All simulation scenario modeling is performed in the open-source agent-based SUMO microscopic traffic simulator (33).
@@
In vehicle-level modeling, we aim to capture the factors owing to the operating characteristics of individual vehicle engines. These factors are modeled using vehicle emission models. This approach allows us to account for factors such as temperature, humidity, seasonality, and weather conditions. The industry-standard Motor Vehicle Emission Simulator (MOVES) by the US Environmental Protection Agency (49) provides accurate exhaust emissions for this purpose. Yet, it is impractical for our large-scale analysis, where we must assess millions of scenarios, given the MOVES takes about three minutes per scenario. To overcome this, we create a suite of efficient surrogate neural network emission models using MOVES as the data source (42; 40). Leveraging MOVES also allows us to capture the cumulative average effects of vehicle maintenance, vehicle model, vehicle characteristics, road types, aerodynamic drag, air conditioning, and vehicle load.
Last, we focus on behavior-level modeling, which accounts for the factors owing to altered vehicle behaviors due to eco-driving optimization. These factors are modeled using deep reinforcement learning and rule-based control modeling. We consider behaviors with a substantial impact on emissions, such as idling, speeding, acceleration, deceleration, pulse-and-glide, and lane changes (61). Behavior-level modeling across diverse traffic scenarios is a core challenge addressed in this work, as discussed next.
@@ 6.2 Eco-driving Control
Eco-driving control involves devising a set of eco-driving strategies for the representative scenario set $\Phi^{{}^{\prime}}$. Eco-driving control for one million scenarios is a challenging task. Traditional methods like dynamic programming are computationally limiting (10), while heuristics are often suboptimal (27). Trajectory optimizations are limited by the complexity of required vehicle dynamics models, particularly due to inter-vehicle dynamics and environmental variability introduced by many factors we consider (64).
Meanwhile, recent studies have shown the effectiveness of deep reinforcement learning in complex control tasks (60; 44; 5). Neural vehicle control using deep reinforcement learning (58; 62; 25), which can handle diverse traffic scenarios without manual fine-tuning, is also gaining popularity. However, training such a policy for each scenario is impractical due to the number of scenarios. Training a unified policy is complex due to multifaceted factors and the induced complex expectation (65). Learning both discrete lane changes and continuous longitudinal accelerations presents optimization challenges, demanding diverse exploration strategies and distinct neural network representations.
@@
Similar scenario partitioning: In this step, our aim is to partition scenarios by their similarities, allowing us to train separate MT-DRL policies for each partition. This strategy is inspired by the mixture of expert architecture widely used in real-world systems with generalization requirements (35). It can reduce scenario diversity within a partition and help reduce training complexity. While each scenario is defined by many factors, we only utilize intersection geometry and traffic flow-specific features as partitioning features because they strongly influence policy behaviors.
However, since signalized intersections lack rotational invariance, comparing them for similarity is challenging. Therefore, we break down the intersection environment into non-conflicting pairs of incoming and outgoing approaches. This has added advantages. It accelerates training by simulating fewer vehicles and improves training stability by alleviating credit assignment challenges (66). Further, it has no impact on modeling realism due to the decoupling nature of traffic signals (e.g., northbound traffic and eastbound traffic do not collide). These selected pairs, along with the ghost approaches connected to them, thus become our training environments and are illustrated in Figure 7.
@@
The training environments are then partitioned into 28 partitions, considering the semi-autonomous vehicle penetration level, lane count, and traffic signal phase count as partition features. However, the environments in each partition may not form a uniform distribution. This may make environments sampled in training biased towards certain types of intersections for the most part (implicit imbalance problem). To overcome this, we adopt a synthetic environment distribution $\Phi_{T}$ derived from the full scenario set $\Phi$ for each partition with a uniform distribution for policy training as illustrated in Figure 2. Additional details about the training environment distributions can be found in Appendix 10.7.
@@
Observation: Vehicle observation design is mainly influenced by the currently available technology, such as the requisite vehicle sensors (e.g., vehicle-to-infrastructure communication range), and the significance attributed by existing literature to specific information required for eco-driving. Thus, we define the observation for a vehicle (agent) $i$ as follows.
@@
Reward: The reward function must encourage algorithmic generalization within the CMDP while considering various competing objectives like emission reduction and travel time efficiency, as we introduced in Section 5.2. Thus, the reward function for each semi-autonomous vehicle is expressed as $r_{t}^{i}=v_{t}^{i}+\alpha{1}_{v^{i}_{t}<\upsilon}+\beta e_{t}^{i}$, where $v_{t}^{i}$ represents the velocity of vehicle $i$ at time $t$, and $e_{t}^{i}$ represents its CO2 emissions. In this equation, $\alpha$, $\beta$, and $\upsilon$ are hyperparameters. The indicator function ${1}_{v^{i}_{t}<\upsilon}$ indicates whether the vehicle is stopped, while the term $e_{t}^{i}$ encourages low emissions. The velocity term captures the effect on travel time. We note that we do not perform any explicit reward shaping to ensure the constraints of the regional eco-driving problem (Eq 8, 9, and 10) are satisfied.
It’s generally accepted that limiting accelerations and idling reduces emissions (38). But this may create traffic overflows or reduce intersection throughput. To account for these observations, we devise two policies per partition ${\psi_{i}}$: one, denoted as $\pi_{\psi_{i}}^{1}$, penalizes idling, while the other, $\pi_{\psi_{i}}^{2}$, places less weight on it. To achieve this we set $\alpha=35$, $\beta=3$ and $\upsilon=1$ for $\pi_{\psi_{i}}^{1}$ and $\alpha=10$, $\beta=3$ and $\upsilon=1$ for $\pi_{\psi_{i}}^{2}$.
Hierarchical controller architecture. While our MT-DRL policy controls vehicle accelerations, we leverage a rule-based default SUMO lane-changing controller for lane changes. It captures 1) strategic lane changes, 2) cooperative lane changes, 3) tactical lane changes, and 4) regulatory lane changes (13). At every time step, we check if the lane-changing controller would predict a lane-changing for the vehicle. If a lane-changing maneuver is warranted, precedence is given to the lane-changing maneuvers, and the lane-changing controller takes over the control of the vehicle to control both longitudinal and lateral movements of the vehicle for that time step. Otherwise, we use the MT-DRL policy by default to control the longitudinal movements of the vehicle.